Comparative financial statements for Weller Corporation, a merchandising company, for the year ending December 31 appear below. The company did not issue any new common stock during the year. A total of 600,000 shares of common stock were outstanding. The interest rate on the bond payable was 10%, the income tax rate was 40%, and the dividend per share of common stock was $0.75 last year and $0.40 this year. The market value of the company’s common stock at the end of this year was $21. All of the company’s sales are on account.

Weller Corporation
Comparative Balance Sheet
(dollars in thousands)
This Year Last Year
Assets
Current assets:
Cash $ 1,080 $ 1,350
Accounts receivable, net 10,300 6,700
Inventory 13,300 12,100
Prepaid expenses 630 630
Total current assets 25,310 20,780
Property and equipment:
Land 10,500 10,500
Buildings and equipment, net 47,698 40,894
Total property and equipment 58,198 51,394
Total assets $ 83,508 $ 72,174
Liabilities and Stockholders' Equity
Current liabilities:
Accounts payable $ 18,600 $ 18,700
Accrued liabilities 1,020 740
Notes payable, short term 0 270
Total current liabilities 19,620 19,710
Long-term liabilities:
Bonds payable 8,600 8,600
Total liabilities 28,220 28,310
Stockholders' equity:
Common stock 2,000 2,000
Additional paid-in capital 4,000 4,000
Total paid-in capital 6,000 6,000
Retained earnings 49,288 37,864
Total stockholders' equity 55,288 43,864
Total liabilities and stockholders' equity $ 83,508 $ 72,174
Weller Corporation
Comparative Income Statement and Reconciliation
(dollars in thousands)
This Year Last Year
Sales $ 72,000 $ 66,000
Cost of goods sold 34,000 35,000
Gross margin 38,000 31,000
Selling and administrative expenses:
Selling expenses 10,800 10,200
Administrative expenses 6,900 6,900
Total selling and administrative expenses 17,700 17,100
Net operating income 20,300 13,900
Interest expense 860 860
Net income before taxes 19,440 13,040
Income taxes 7,776 5,216
Net income 11,664 7,824
Dividends to common stockholders 240 600
Net income added to retained earnings 11,424 7,224
Beginning retained earnings 37,864 30,640
Ending retained earnings $ 49,288 $ 37,864

Required:

Compute the following financial ratios for this year:

1. Times interest earned ratio.

2. Debt-to-equity ratio.

3. Equity multiplier.
Answer #1
1. Times interest earned ratio
Net operating income before   / Interest expense = Times interest earned ratio
interest expense and income taxes
                                                   20,300.00 /                                                   860.00 =              23.60
2. Debt to equity ratio
Total liabilities / Stockholders’ equity = Debt to equity ratio
                                                   28,220.00 /                                             55,288.00 =                 0.51 :1
3. Equity multiplier
Average total assets   / Average stock holder’s equity = Equity multiplier
                                                   77,841.00 /                                             49,576.00 =                 1.57
Explanation
Average total assets   = (83,508 + 72,174)/2 = $ 77,841.00
Average stock holder’s equity = (55,288 + 43,864)/2 = $ 49,576.00